In a collision of sufficient force, automobile airbags respond by electrically triggering the explosive decomposition of sodium azide (NaN3) to its elements. A 29.4−g sample of sodium azide was... decomposed, and the nitrogen gas generated was collected over water at 26°C. The total pressure was 733.8 mmHg. How many liters of dry N2 were generated?


1
Expert's answer
2021-02-28T06:19:11-0500

Molar mass of NaN3 = 65g

65g of NaN3 yields 28g of N2

29.4g of NaN3 yields (29.4x28)/65= 12.66g of N2

Amount of N2 produced by 29.4g of NaN3 = 12.66/28 = 0.45 mol

Pressure= 733.8/760 = 0.9655atm

Temperature= 26°C = 299K

R= 8.314 J^-1K^-1mol^-1


Using PV = nRT

V= nRT/P

= (0.45x8.314x299)/0.9655

=1158.62L